You measure 30 textbooks' weights, and find they have a mean weight of 78 ounces. Assume the population standard deviation is 2.3 ounces. Based on this, construct a 99% confidence interval for the true population mean textbook weight.
Give your answers as decimals, to two places
< μ <

Answer by math_tutor2020(3816)


Given information:
mean = xbar = 78
standard deviation = sigma = 2.3
sample size = n = 30
confidence level = 99%

A 99% confidence level means the critical value is approximately z = 2.58 which you find using a calculator or table.

The lower value of the confidence interval is
L = xbar - z*sigma/sqrt(n)
L = 78 - 2.58*2.3/sqrt(30)
L = 78 - 1.083395
L = 76.916605
L = 76.92

The upper value of the confidence interval is
U = xbar + z*sigma/sqrt(n)
U = 78 + 2.58*2.3/sqrt(30)
U = 78 + 1.083395
U = 79.083395
U = 79.08

The 99% confidence interval is therefore
76.92 < μ < 79.08